OpenStudy (anonymous):

6. Cleo and Clare are looking from their balcony to a swimming pool below that is located 15 m horizontally from the bottom of their building. They estimate the balcony is 45 m high and wonder how fast they would have to jump horizontally to succeed in reaching the pool. What calculations would you show to help them determine the answer? Evaluate the practicality of their being able to succeed at jumping into the pool.

OpenStudy (jack1):



OpenStudy (jack1):

g = 9.8 m/s^2 s = ut + 1/2 a t^2 \(\large s = ut + \frac12 a t^2\) where u = 0 a = g = 9.8 t = ? s = 45 solve for t

OpenStudy (jack1):

\(\large s = ut + \frac12 a t^2\) \(\large 45 = 0 + \frac12 g t^2\) \(\large 45 = 4.9 t^2\) \(\large 9.184 = t^2\) t = 3.03 seconds from balcony to splat

OpenStudy (jack1):

so, to safely make it into the pool, cleo/clare need to travel 16 m in 3.03 seconds = horizontal speed of 5.28 m/s thats a jog, not a sprint, so plenty safe

OpenStudy (jack1):

the entry speed into the water is a bit fast tho... 3.03 x 9.8 roughly = 30m/s = 107 kms/hr into the water.. so probably like landing on concrete at that speed

OpenStudy (mrnood):

BEWARE!!! Even professional cliff divers are diving from a height of approx. 30m 45m is WAY too high for anyone ot safely attempt to jump from it!

OpenStudy (jack1):

maybe if they jumped with some rocks... to break the surface tension... then it'd be ok

OpenStudy (mrnood):

@Jack1 surface tension is NOT relevant to this question - it acts as a surface force, horizontally only. Rocks will not help The key point here is that the body moving through water has massive 'drag' due to the density of the water compared to air. Therefore the acceleration upon hitting the water is extremely high and is liable to cause internal injuries from the 'shock'.

OpenStudy (mrnood):

If we model the body jumping as a cylinder with cross section of say 400mm diameter the initial force from entering the water is I/2 rho v^2 A (rho = density = 1000kg/m^3) = 500 * 30^2 *pi * 0.2^2 54kN If it a big jumper of mass 100kg then the acceleration = 54000/100 = 540 m/s^2 i.e. 50g Bodies cannot normally withstand this type of acceleration without injury.

OpenStudy (jack1):

@MrNood incorrect, breaking the surface tension allows for greater entry speed and creates micro-currents travelling in the same direction as the diver, allowing for a slower de-acceleration

OpenStudy (mrnood):

it is a totally insignificant effect compared to the drag force

OpenStudy (mrnood):

and in any case - the gross movement of the water caused by the rock is not related to surface tension - it simply reduces the effective velocity of entry. Surface tension is a horizontal 'skin' effect and has no relevance to a body breaking into the fluid in macro scale

OpenStudy (jack1):